OpenStudy (anonymous):

The ratio between the radii of two sphere is 9.2. What is the radii of their volume?

Directrix (directrix):

If two solids are similar, the cube of the scale factor of the two solids is equal to the ratio of the volumes.

Directrix (directrix):

>>What is the radii of their volume? Did you mean to write: What is the ratio of their volumes?

OpenStudy (anonymous):

Yes

Directrix (directrix):

Okay.

Directrix (directrix):

The ratio of the radii is the scale factor. It is given to be 9:2 or 9/2. Your task is this: What is the cube of (9/2) ? (9/2)^3 = ? That will be the ratio of the volumes. Post what you get, okay? @reymishia

OpenStudy (anonymous):

27

OpenStudy (anonymous):

And it's giving to me like this 9:2

Directrix (directrix):

No. Try again. 9 * 9 * 9 --------- = ? 2 * 2* 2

Directrix (directrix):

What is 9 times 9 times 9 equal?

OpenStudy (anonymous):

729

Directrix (directrix):

What is 2 * 2 * 2 =

OpenStudy (anonymous):

8

Directrix (directrix):

The ratio of the volumes is the fraction: 729 ---- 8 You can also write it as 729 : 8

OpenStudy (anonymous):

OK thanks

Directrix (directrix):

You are welcome.